# Breaking the Exploration Bottleneck: Rubric-Scaffolded Reinforcement Learning for General LLM Reasoning

### v5

@@ 6 Conclusion and Discussion
In this work, we apply instructional scaffolding theory from educational psychology to RL for LLMs, and introduce RuscaRL, a novel instructional scaffolding framework that breaks the exploration bottleneck for general LLM reasoning tasks. RuscaRL leverages checklist-style rubrics through scaffolding mechanisms that provide gradually decaying external guidance and reward functions that enable robust RL training. The rubric-based scaffolding mechanism provides external guidance that gradually decays to encourage internalization, while the rubric-based reward function enables robust evaluation for effective RL training. Extensive experiments demonstrate that RuscaRL consistently outperforms strong baseline methods and achieves competitive results compared with leading models. Furthermore, the model fine-tuned with RuscaRL produces highly novel responses that the initial model could barely generate.
While RuscaRL demonstrates promising results in breaking the exploration bottleneck for general LLM reasoning, several limitations remain that highlight directions for future research. Our approach critically relies on high-quality, well-structured rubric datasets, which are still scarce in the community, and is highly sensitive to rubric design quality. Poorly designed rubrics may fail to provide robust reward signals due to unreasonable point allocations or conflicting criteria, while narrow rubrics may limit the scaffolding process’s ability to generate diverse, high-quality responses. The success of RuscaRL underscores the urgent need for broader community investment in constructing open, diverse, and domain-rich rubric datasets. Our future work includes developing pipelines for high-quality rubric data production, exploring rubric-based natural language feedback strategies, and investigating applications to multi-modal tasks and agent systems.
Appendix
@@ Appendix A Additional Related Works
LLM Reasoning. While early methods like prompt engineering (Wei et al., 2022; Kojima et al., 2022) and supervised fine-tuning (Ouyang et al., 2022) have yielded encouraging results, their reliance on task-specific prompts or extensive labeled data limits their scalability and cross-domain generalization (Stiennon et al., 2020; Pornprasit & Tantithamthavorn, 2024; Zhang et al., 2024b; Gao et al., 2023). Recent works have found that using more test-time computation (Snell et al., 2024; Zhang et al., 2024a; Zuo et al., 2025) can improve the reasoning performance of LLMs. More recently, RLVR (Lambert et al., 2024; OpenAI, 2025a; Guo et al., 2025) has emerged as a promising paradigm for training LLMs to solve verifiable problems, showing strong reasoning improvements in domains like math and coding (Guo et al., 2025; Qwen, 2025; Lambert et al., 2024; OpenAI, 2025a). However, it faces a significant exploration bottleneck (Wu et al., 2025a; Yue et al., 2025; Liu et al., 2025b) and is difficult to extend to general tasks where correctness is hard to verify (Gunjal et al., 2025; Kimi et al., 2025).
@@ Appendix B Algorithm Pseudocode
5:   Compute scaffolding ratio vector: $\bm{\lambda}_{S}=\lambda_{step}(t)\times\bm{\lambda}_{group}$
@@ C.1 Detailed Training Settings
Training Datasets. For the medical domain, we use the remaining 4500 samples from HealthBench after excluding HealthBench-500. For the other domains, we generate HealthBench-like rubrics data by calling GPT-4.1 (OpenAI, 2025b) with specific prompts detailed in Appendix E.3. For the writing domain, we combine LongWriter-6k (Bai et al., 2024) and LongWriter-Zero-RLData (Wu et al., 2025b) datasets. For the instruction following domain, we use IF-multi-constraints-upto5 (Pyatkin et al., 2025) dataset. For the STEM domain, we use SCP-116K (Lu et al., 2025) and MATH training datasets (Level 3-5) (Hendrycks et al., 2021).
Training Configurations. This section provides detailed training configurations, as shown in Table 3. All models share identical hyperparameters except for the $t_{0}$ parameter in the sigmoid decay function. Specifically, Qwen3-30B-A3B-Instruct and Qwen3-30B-A3B-Base use $t_{0}=0.1$, Llama-3.1-8B-Instruct and Llama-3.1-8B use $t_{0}=0.15$, Llama-3.2-3B-Instruct uses $t_{0}=0.3$, and the remaining models (Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-7B, Qwen2.5-32B-Instruct, Qwen2.5-32B, Qwen3-4B-Instruct-2507 and Qwen3-4B-Base) use $t_{0}=0.2$.
@@ C.2 Detailed Evaluation Settings
For medical benchmarks (HealthBench-500 and LLMEval-Med), we employ GPT-4.1 as the judge model. For writing benchmarks (WritingBench and Creative Writing v3), we employ Claude-Sonnet-4 as the judge model. Our generation parameters are set to Temperature=0.7, Top-P=0.8, and Top-K=20 across all evaluations. The maximum output length is configured as 4096 tokens for non-writing tasks and 16000 tokens for writing tasks. For IFEVAL and IFBench, we report the prompt-level strict-accuracy metric. We report single evaluation results for HealthBench-500, LLMEval-Med, and WritingBench, while for MedQA, MedMCQA, Creative Writing v3, IFEVAL, IFBench, GPQA-D, MMLU, MMLU-Pro, MATH-500, AMC 2023, AIME 2024, and AIME 2025, we report the average of three runs. All scores are converted to a percentage scale for reporting.
We also compare against other models, including closed-source models (OpenAI-o3 (OpenAI, 2025a), GPT-4.1 (OpenAI, 2025b), Gemini-2.5-Pro (Google, 2025)) and open-source models (DeepSeek-R1-0528 (Guo et al., 2025), Qwen3-235B-Thinking-2507 (Yang et al., 2025), Kimi-K2-Instruct (Kimi et al., 2025), gpt-oss-120b, gpt-oss-20b (OpenAI, 2025c), Rubicon-Preview (Huang et al., 2025)), on HealthBench-500 (Figure 1), to demonstrate the competitiveness of our approach.
Appendix D Detailed Experimental Analysis
@@ D.1 Performance Across Different Models
Llama-3.1-8B-Instruct 12.4 46.0 29.8 46.3 66.8 70.7 58.0 60.7
Llama-3.1-8B 0 25.8 9.1 29.5 36.9 49.7 35.9 45.4
Qwen2.5-7B 8.2 46.2 28.2 47.8 55.3 58.2 55.0 55.6
@@
Qwen3-30B-A3B-Base 11.4 48.3 43.3 60.9 73.6 71.3 65.1 65.4
D.2 Mixed Training Analysis
To evaluate the effectiveness of different training strategies, we compare domain-specific training, health-only training, and mixed training approaches on Qwen2.5-7B-Instruct. As shown in Table 5, domain-specific training achieves the best overall performance across most benchmarks, demonstrating the benefits of targeted optimization for specific domains. Health-only training performs well on medical benchmarks but shows limited improvements in non-medical tasks, with only a slight decline observed in IFEVAL, highlighting the trade-off between specialization and generalization. Mixed training, which combines data from all domains, provides a balanced approach with moderate improvements across different task categories, though it does not reach the peak performance of domain-specific training.
Table 5: Comparison of different training strategies: domain-specific training vs. health-only training vs. mixed training on Qwen2.5-7B-Instruct.
@@ D.2 Mixed Training Analysis
Initial 23.6 47.9 61.8 56.3 45.3 37.4 71.0 26.8
Domain-specific Training 50.3 (+26.7) 61.2 (+13.3) 63.5 (+1.7) 56.5 (+0.2) 56.4 (+11.1) 38.6 (+1.2) 75.3 (+4.3) 31.0 (+4.2)
Health-only Training 50.3 (+26.7) 61.2 (+13.3) 63.5 (+1.7) 56.5 (+0.2) 55.8 (+10.5) 35.1 (-2.3) 68.0 (-3.0) 27.2 (+0.4)
Mixed Training 44.3 (+20.7) 56.7 (+8.8) 62.7 (+0.9) 56.8 (+0.5) 50.4 (+5.1) 35.6 (-1.8) 71.2 (+0.2) 33.7 (+6.9)
D.3 Supervised Fine-tuning vs. RuscaRL
As shown in Table 6, SFT using GPT-4.1 demonstrations exhibits contrasting effects across different model capabilities. For weaker models like Qwen2.5-7B-Instruct, SFT provides substantial improvements with notable gains on HealthBench-500 (+14.7) and WritingBench (+17.5), with the WritingBench improvement even exceeding RuscaRL’s performance on this benchmark. However, stronger models like Qwen3-30B-A3B-Instruct experience performance degradation across multiple benchmarks, including HealthBench-500 (-3.0), and WritingBench (-12.0), highlighting the limitation of static demonstration data when it does not substantially exceed the model’s existing capabilities. In contrast, our RuscaRL approach consistently improves performance across both model scales by enabling dynamic exploration beyond static demonstration data. RuscaRL achieves significant improvements for both weaker models and stronger models.
@@
Initial 46.8 71.4 84.2 71.3 78.4 74.4 83.0 31.9
SFT 43.8 (-3.0) 65.7 (-5.7) 82.0 (-2.2) 70.3 (-1.0) 66.4 (-12.0) 62.7 (-11.7) 83.1 (+0.1) 30.2 (-1.7)
RuscaRL 61.1 (+14.3) 73.0 (+1.6) 84.8 (+0.6) 71.9 (+0.6) 79.2 (+0.8) 74.3 (-0.1) 84.5 (+1.5) 32.1 (+0.2)
D.4 Additional Metrics Analysis
@@
D.4.1 Novelty Analysis
@@
To validate that RuscaRL achieves significantly higher novelty improvement compared to Rubric-based RL after training. Table 7 shows the performance of both methods in terms of importance ratios. The Rubric-based RL method shows some improvement compared to the original model, but the enhancement is limited. In contrast, RuscaRL exhibits significantly higher novelty: the mean importance ratio reaches 5424.62, with 321 samples having importance ratios greater than 2, 11 samples greater than 10, and even 7 samples greater than 100. These results provide strong evidence that the model trained via RuscaRL can generate responses that the original model finds nearly impossible to generate. As shown in Figure 7, RuscaRL demonstrates clear advantages in novelty metrics.
@@
RuscaRL demonstrates significantly higher importance ratios than Rubric-based RL, with the top sample reaching $\rho_{seq}=2,638,481.94$ compared to Rubric-based RL’s maximum of $35.66$, indicating more aggressive policy space exploration. Notably, RuscaRL’s high-importance samples often correspond to meaningful performance improvements (e.g., score differences of $0.54$, $0.89$, $0.67$, $0.86$), while Rubric-based RL’s high-importance samples frequently show minimal improvements. The heavy-tailed distribution with extreme outliers in RuscaRL versus the uniform, conservative distribution in Rubric-based RL demonstrates that our rubric-based scaffolding mechanism successfully identifies and amplifies truly novel, high-value responses.
@@
D.4.2 Diversity Analysis
@@ D.5 Importance Sampling Analysis
Initial 23.6 47.9 61.8 56.3
$\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}$ 50.3 61.2 63.5 56.5
@@
D.6 Training Runtime
The training process consists of three stages: Rollout, Reward, and Actor Update. Notably, RuscaRL maintains a similar training runtime as rubric-based RL methods (Gunjal et al., 2025; Huang et al., 2025). For Qwen2.5-7B-Instruct on the medical task using 8 × H200 GPUs, the average time per step is 40 seconds for Rollout and 15 seconds for Actor Update. Reward time depends on service efficiency (e.g., vLLM (Kwon et al., 2023), SGLang (Zheng et al., 2024)) or external API calls. In our experiments using external APIs, the average Reward time per step is 20 seconds. Therefore, training for 5 epochs (350 steps) requires a total of 7.3 hours.

### v6

@@ 6 Conclusion and Discussion
In this work, we apply instructional scaffolding theory to RL for LLMs, and introduce RuscaRL, a novel instructional scaffolding framework that breaks the exploration bottleneck for general LLM reasoning tasks. RuscaRL leverages checklist-style rubrics through scaffolding mechanisms that provide gradually decaying external guidance to encourage internalization, while the rubric-based reward function enables robust and fine-grained evaluation for effective RL training on general tasks. Extensive experiments demonstrate that RuscaRL consistently outperforms strong baseline methods and achieves competitive results compared with leading models. Furthermore, the model fine-tuned with RuscaRL produces highly novel responses that the initial model could barely generate, effectively expanding the reasoning boundary. The success of RuscaRL underscores the urgent need for broader community investment in constructing open, diverse, and domain-rich rubric datasets. Our future work includes developing pipelines for high-quality rubric data production, exploring rubric-based natural language feedback strategies, and investigating applications to multi-modal tasks and agent systems.
@@ Appendix A Additional Related Works
LLM Reasoning. While early methods like prompt engineering (Wei et al., 2022; Kojima et al., 2022; Zhou et al., 2023a; Yao et al., 2023b, a) and supervised fine-tuning (Ouyang et al., 2022) have yielded encouraging results, their reliance on task-specific prompts or extensive labeled data limits their scalability and cross-domain generalization (Stiennon et al., 2020; Pornprasit and Tantithamthavorn, 2024; Zhang et al., 2024b; Gao et al., 2023). Recent works have found that using more test-time computation (Snell et al., 2024; Zhang et al., 2024a; Zuo et al., 2025) can improve the reasoning performance of LLMs. More recently, RLVR (Lambert et al., 2024; OpenAI, 2025c; Guo et al., 2025) has emerged as a promising paradigm for training LLMs to solve verifiable problems, showing strong reasoning improvements in domains like math and coding (Guo et al., 2025; Qwen, 2025; Lambert et al., 2024; OpenAI, 2025c). However, it faces a significant exploration bottleneck (Wu et al., 2025a; Yue et al., 2025; Liu et al., 2025c) and is difficult to extend to general tasks where correctness is hard to verify (Gunjal et al., 2025; Kimi et al., 2025).
@@ Appendix B Algorithm Pseudocode
5:   Compute scaffolding ratio vector: $\boldsymbol{\lambda}_{S}=\lambda_{step}(t)\times\boldsymbol{\lambda}_{group}$
@@ C.1 Detailed Training Settings
Training Datasets. For the medical domain, we use the remaining 4500 samples from HealthBench after excluding HealthBench-500. For the other domains, we generate HealthBench-like rubrics data by calling GPT-4.1 (OpenAI, 2025a) with specific prompts detailed in Appendix E.3. For the writing domain, we combine LongWriter-6k (Bai et al., 2024) and LongWriter-Zero-RLData (Wu et al., 2025b) datasets. For the instruction following domain, we use IF-multi-constraints-upto5 (Pyatkin et al., 2025) dataset. For the STEM domain, we use SCP-116K (Lu et al., 2025) and MATH training datasets (Level 3-5) (Hendrycks et al., 2021).
Training Configurations. This section provides detailed training configurations, as shown in Table 3. All models share identical hyperparameters except for the $t_{0}$ parameter in the sigmoid decay function and the sampling parameters. For the $t_{0}$ parameter, Qwen3-30B-A3B-Instruct and Qwen3-30B-A3B-Base use $t_{0}=0.1$, Llama-3.1-8B-Instruct and Llama-3.1-8B use $t_{0}=0.15$, Llama-3.2-3B-Instruct uses $t_{0}=0.3$, and the remaining models (Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-7B, Qwen2.5-32B-Instruct, Qwen2.5-32B, Qwen3-4B-Instruct-2507 and Qwen3-4B-Base) use $t_{0}=0.2$. For sampling parameters, we use a conservative decoding strategy (Temperature=0.7, Top-P=0.8, Top-K=20) for most models. However, we observed that Qwen2.5-7B-Instruct exhibits rapid policy entropy collapse under this configuration. Therefore, for Qwen2.5-7B-Instruct, we use Temperature=1.0 with Top-P and Top-K disabled.
@@ C.2 Detailed Evaluation Settings
For medical benchmarks (HealthBench-500 and LLMEval-Med), we employ GPT-4.1 as the judge model. For writing benchmarks (WritingBench and Creative Writing v3), we employ Claude-Sonnet-4 as the judge model. Our generation parameters are set to Temperature=0.7, Top-P=0.8, and Top-K=20 across all evaluations. The maximum output length is configured as 4096 tokens for non-writing tasks and 16000 tokens for writing tasks. For IFEVAL and IFBench, we report the prompt-level strict-accuracy metric. We report the average of three runs for all evaluation benchmarks. All scores are converted to a percentage scale for reporting.
We also compare against other models, including closed-source models (OpenAI-o3 (OpenAI, 2025c), GPT-4.1 (OpenAI, 2025a), Gemini-2.5-Pro (Google, 2025)) and open-source models (DeepSeek-R1-0528 (Guo et al., 2025), Qwen3-235B-Thinking-2507 (Yang et al., 2025), Kimi-K2-Instruct (Kimi et al., 2025), gpt-oss-120b, gpt-oss-20b (OpenAI, 2025b), Rubicon-Preview (Huang et al., 2025)), on HealthBench-500 (Figure 1), to demonstrate the competitiveness of our approach.
Appendix D Additional Experimental Analysis
@@ D.1 Performance Across Different Models
Llama-3.1-8B-Instruct 12.5 46.0 30.1 46.2 66.8 70.7 58.0 60.7
Llama-3.1-8B 0.0 25.8 9.1 29.6 36.9 49.7 35.9 45.4
Qwen2.5-7B 8.5 46.3 28.2 47.9 55.3 58.2 55.0 55.6
@@
Qwen3-30B-A3B-Base 11.2 48.4 43.1 60.9 73.6 71.3 65.1 65.4
D.2 Performance in the STEM Domain
Figure 5: STEM Benchmarks.
For the STEM domain, we use GPQA-Diamond (Rein et al., 2024), MMLU (Hendrycks et al., 2020), MMLU-Pro (Wang et al., 2024), MATH-500 (Lightman et al., 2023), AMC 2023Available at https://huggingface.co/datasets/knoveleng/AMC-23., AIME 2024Available at https://huggingface.co/datasets/HuggingFaceH4/aime_2024., and AIME 2025Available at https://huggingface.co/datasets/opencompass/AIME2025.. Detailed evaluation settings are provided in Appendix C.2. Meanwhile, RuscaRL has also been successfully extended to the STEM domain: experiments on Qwen2.5-7B-Instruct show consistent performance improvements across all STEM benchmarks (see Figure 5).
D.3 Mixed Training Analysis
To evaluate the effectiveness of different training strategies, we compare domain-specific training, health-only training, and mixed training approaches on Qwen2.5-7B-Instruct. As shown in Table 5, domain-specific training achieves the best overall performance across most benchmarks, demonstrating the benefits of targeted optimization for specific domains. Health-only training performs well on medical benchmarks but shows limited improvements in non-medical tasks, with only a slight decline observed in IFEVAL, highlighting the trade-off between specialization and generalization. Note that health-only training uses the same medical training data as domain-specific training, resulting in identical performance on medical benchmarks. Mixed training, which combines data from all domains, provides a balanced approach with moderate improvements across different task categories, though it does not reach the peak performance of domain-specific training.
Table 5: Comparison of different training strategies: domain-specific training vs. health-only training vs. mixed training.
@@ D.3 Mixed Training Analysis
Initial 23.4 48.0 61.8 56.3 45.2 37.4 71.0 26.8
Domain-specific Training 56.4 (+33.0) 65.3 (+17.3) 63.5 (+1.7) 56.5 (+0.2) 56.1 (+10.9) 38.6 (+1.2) 75.3 (+4.3) 31.0 (+4.2)
Health-only Training 56.4 (+33.0) 65.3 (+17.3) 63.5 (+1.7) 56.5 (+0.2) 55.8 (+10.6) 35.1 (-2.3) 68.0 (-3.0) 27.2 (+0.4)
Mixed Training 54.2 (+30.8) 58.3 (+10.3) 62.5 (+0.7) 56.3 (+0.0) 50.7 (+5.5) 35.7 (-1.7) 71.2 (+0.2) 33.8 (+7.0)
D.4 Supervised Fine-tuning vs. RuscaRL
As shown in Table 6, SFT using GPT-4.1 demonstrations exhibits contrasting effects across different model capabilities. For weaker models like Qwen2.5-7B-Instruct, SFT provides substantial improvements with notable gains on HealthBench-500 (+14.9) and WritingBench (+17.6), with the WritingBench improvement even exceeding RuscaRL’s performance on this benchmark. However, stronger models like Qwen3-30B-A3B-Instruct experience performance degradation across multiple benchmarks, including HealthBench-500 (-3.1), and WritingBench (-11.7), highlighting the limitation of static demonstration data when it does not substantially exceed the model’s existing capabilities. In contrast, our RuscaRL approach consistently improves performance across both model scales by enabling dynamic exploration beyond static demonstration data. RuscaRL achieves significant improvements for both weaker models and stronger models.
@@
Initial 46.9 71.5 84.2 71.3 78.1 74.4 83.0 31.9
SFT 43.8 (-3.1) 65.7 (-5.8) 82.0 (-2.2) 70.3 (-1.0) 66.4 (-11.7) 62.7 (-11.7) 83.1 (+0.1) 30.2 (-1.7)
RuscaRL 61.1 (+14.2) 73.2 (+1.7) 84.8 (+0.6) 71.9 (+0.6) 79.2 (+1.1) 74.3 (-0.1) 84.5 (+1.5) 32.1 (+0.2)
D.5 Additional Metrics Analysis
@@
D.5.1 Novelty Analysis
To validate that RuscaRL achieves substantially higher novelty improvement compared to Rubric-based RL after training. Table 7 shows the performance of both methods in terms of importance ratios. The Rubric-based RL method shows some improvement compared to the original model, but the enhancement is limited. In contrast, RuscaRL exhibits substantially higher novelty: the mean importance ratio reaches 5424.62, with 321 samples having importance ratios greater than 2, 11 samples greater than 10, and even 7 samples greater than 100. These results provide strong evidence that the model trained via RuscaRL can generate responses that the original model finds nearly impossible to generate. As shown in Figure 6, RuscaRL demonstrates clear advantages in novelty metrics.
@@ D.5.1 Novelty Analysis
RuscaRL demonstrates substantially higher importance ratios than Rubric-based RL, with the top sample reaching $\rho_{seq}=2,638,481.94$ compared to Rubric-based RL’s maximum of $35.66$, indicating more aggressive policy space exploration. Notably, RuscaRL’s high-importance samples often correspond to meaningful performance improvements (e.g., score differences of $0.54$, $0.89$, $0.67$, $0.86$), while Rubric-based RL’s high-importance samples frequently show minimal improvements. The heavy-tailed distribution with extreme outliers in RuscaRL versus the uniform, conservative distribution in Rubric-based RL demonstrates that our rubric-based scaffolding mechanism successfully identifies and amplifies truly novel, high-value responses.
@@
D.5.2 Diversity Analysis
D.6 Importance Sampling Analysis
@@
Initial 23.4 48.0 61.8 56.3
$\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}$ 56.4 65.3 63.5 56.5
@@
D.7 Training Runtime
The training process consists of three stages: Rollout, Reward, and Actor Update. Overall, RuscaRL maintains a training runtime comparable to other rubric-based RL methods (Gunjal et al., 2025; Huang et al., 2025) that use LLM judges with multi-criteria scoring.
In our main experiments on the medical task, the policy model (e.g., Qwen2.5-7B-Instruct) is trained on one $8\timesH$200 node, and the Grader model (Qwen3-32B, non-thinking) on an additional node. For each step, we use a batch size of 64 instructions, 8 rollouts per instruction, and an average of 11.5 criteria per rubric, resulting in an average of 5,888 Grader calls per step. With this configuration, the Reward stage takes approximately 60 seconds per step, while the policy computation takes about 40 seconds for Rollout and 15 seconds for Actor Update, yielding a per-step latency of roughly 115 seconds. Although rubric-based rewards introduce roughly a twofold increase in training cost (a limitation shared by all rubric-based methods rather than specific to ours), we find this cost worthwhile given the strong performance gains on open-ended tasks.
Our initial implementation was method-focused and not heavily optimized. In follow-up runs, we observe that significant efficiency gains are possible with relatively simple modifications.
Lightweight grader models. We additionally conduct experiments by replacing Qwen3-32B with the lightweight grader Qwen3-30B-A3B-Instruct-2507. As summarized in Table 10, this modification reduces the per-step Reward-stage wall-clock time from 60 seconds to 18 seconds, with only a slight degradation in final performance on HealthBench-500.
Table 10: Training cost and HealthBench-500 performance with different grader models.
Grader Reward time per training step (s) HealthBench-500 score
Qwen3-32B (non-thinking) 60 56.4
Qwen3-30B-A3B-Instruct-2507 18 48.9
Asynchronous rollout-reward strategy. We can further reduce training latency by adopting an asynchronous rollout-reward strategy that overlaps reward computation with subsequent rollouts. In the default synchronous pipeline, the per-step latency is
$T_{\text{sync}}=T_{\text{rollout}}+T_{\text{reward}}+T_{\text{update}}=40+60+15=115~\text{s/step}.$ (8)
When the Reward stage is run asynchronously, each generated sequence is sent to the grader immediately, and grading is overlapped with subsequent rollouts, the per-step latency becomes
$T_{\text{async}}=\max(T_{\text{rollout}},T_{\text{reward}})+T_{\text{update}}.$ (9)
Under the same configuration as above, this reduces the wall-clock time to
$T_{\text{async}}=\max(40,60)+15=75~\text{s/step}.$ (10)
With the more efficient Qwen3-30B-A3B grader ($T_{\text{reward}}=18~\text{s}$), the latency further drops to
$T_{\text{async}}=\max(40,18)+15=55~\text{s/step},$ (11)
showing that the latency of the Reward stage can be substantially reduced with lightweight graders and simple pipeline optimizations.
D.8 Robustness to Noisy Rubrics
We further quantify its robustness to noisy rubrics. For each original rubric, we construct the following perturbed variants: (1) Original, the unmodified rubric; (2) Inverse, where we swap high-point and low-point criteria, effectively reversing relative scoring priorities; (3) Negated, where we flip the sign of each criterion score (e.g., $3\rightarrow-3$), so that “good” behavior is penalized and “bad” behavior is rewarded; (4) Ambiguous, where we inject vague or subjective criteria generated by GPT-4.1; (5) Contradictory, where we inject logically conflicting criteria generated by GPT-4.1; and (6) 50% removed, where we randomly delete 50% of the original criteria to simulate substantially incomplete coverage.
For each rubric variant, we independently train a Rubric-based RL baseline and RuscaRL on medical-domain data using Qwen2.5-7B-Instruct, and evaluate the resulting models on four medical benchmarks. As shown in Table 11, RuscaRL is consistently more robust to rubric noise: under mild perturbations (Ambiguous, Contradictory, 50% removed), it clearly outperforms Rubric-based RL, while under severe corruptions (Inverse, Negated) both methods degrade substantially.
Table 11: Robustness to rubric noise on medical benchmarks using Qwen2.5-7B-Instruct. Means and standard deviations are computed over three runs.
Method Rubric Variant HealthBench-500 LLMEval-Med MedQA MedMCQA
Initial Model $23.4\pm 0.3$ $48.0\pm 0.3$ $61.8\pm 0.2$ $56.3\pm 0.1$
Rubric-based RL Original $52.0\pm 0.1$ $56.5\pm 0.2$ $62.3\pm 0.4$ $56.3\pm 0.1$
Inverse $8.4\pm 0.3$ $42.6\pm 0.4$ $61.8\pm 0.2$ $56.0\pm 0.2$
Negated $4.3\pm 0.4$ $38.2\pm 1.2$ $61.2\pm 0.1$ $55.8\pm 0.1$
Ambiguous $43.7\pm 0.9$ $55.6\pm 1.0$ $62.1\pm 0.2$ $56.1\pm 0.2$
Contradictory $46.8\pm 0.6$ $56.2\pm 0.9$ $62.0\pm 0.4$ $56.0\pm 0.2$
50% removed $43.2\pm 1.3$ $54.8\pm 0.8$ $61.9\pm 0.5$ $56.2\pm 0.2$
RuscaRL (Ours) Original $\mathbf{56.4}\pm 0.4$ $\mathbf{65.3}\pm 0.5$ $\mathbf{63.5}\pm 0.1$ $\mathbf{56.5}\pm 0.1$
Inverse $11.9\pm 0.6$ $45.7\pm 0.4$ $62.1\pm 0.3$ $56.2\pm 0.2$
Negated $8.2\pm 0.4$ $42.8\pm 0.6$ $61.5\pm 0.1$ $56.0\pm 0.1$
Ambiguous $49.8\pm 0.4$ $62.4\pm 1.2$ $63.1\pm 0.3$ $56.3\pm 0.2$
Contradictory $49.3\pm 0.6$ $60.5\pm 0.7$ $63.2\pm 0.4$ $56.4\pm 0.2$
50% removed $48.1\pm 0.3$ $58.6\pm 0.5$ $63.0\pm 0.3$ $56.2\pm 0.1$